Solve the differential equation y ′′ =1+(y')^2
1
Expert's answer
2020-09-13T18:28:20-0400

Let us substitute "z=y' \Rightarrow y''=z', \; z'=1+z^2."

Therefore, "\dfrac{dz} {1+z^2} =dx."

We should integrate it with respect to x,

"\int\dfrac{dz} {1+z^2} =\int dx, \\
\tan^{-1}z=x+C_1, \\
z=\tan(x+C_1)."

Next, we solve an equation

"y'=\tan (x+C_1), \\
y=\int \tan (x+C_1)dx = - \ln(\cos(x+C_1)) +C_2."